OpenStudy (anonymous):

The length of a screw produced by a machine is normally distributed with a mean of 0.25 inches and a standard deviation of 0.01 inches. What percent of screws are between 0.24 and 0.26 inches? 68% 95% 99.7% 100%

OpenStudy (anonymous):

@maimigirl

OpenStudy (anonymous):

\[\begin{align*}P(0.24<X<0.26)&=P\left(\frac{0.24-0.25}{0.01}<\frac{X-0.25}{0.01}<\frac{0.26-0.25}{0.01}\right)\\\\ &=P(-1<Z<1) \end{align*}\] How much of a population lies within one standard deviation from the mean?

OpenStudy (anonymous):

i was doing such type of questions very recently. population that lies within one standard deviation from the mean is 68%